# Fabian Mauthe

### v2

@@ 18 Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
@@
Signals: position, speed, torque Fault detection Link
@@ 19 Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi Production system
@@
Signals: unknown Fault detection Link
@@ 20 Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@
Signals: flow rate, pressure Prognosis Link
@@ 21 Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS Production system
@@
Signals: anonymized Fault detection Link
@@ 22 Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@
Signals: dust feed, flow rate, pressure Prognosis Link
@@ 23 Kaggle - Pump
Kaggle User: UnknownClass Process technology
@@
Signals: anonymized Fault detection Link
@@ 24 Kaggle - Sensor Fault Detection
Schneider-Electric, France Electrical component
@@
Signals: temperature, voltage Fault detection Link
@@ 25 Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
@@
Signals: control parameters Fault detection Link
@@ 26 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 27 Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2 Drive technology
@@
Signals: current, position, temperature Prognosis Link
@@
Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1 Electrical component
@@
Signals: capacity Prognosis Link
@@ 29 Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil Drive technology
@@
Signals: pressure, temperature, vibration Diagnosis Link
@@ 30 Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3 Mechanical component
@@
Signals: vibration Diagnosis Link
@@ 31 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1 Electrical component
@@
Signals: unknown Prognosis Link
@@ 32 Mendeley - NMC cell 2600 mAh cyclic aging data
Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1 Electrical component
@@
Signals: current, temperature Prognosis Link
@@ 33 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024). Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435. Mechanical component
@@
Signals: vibration Prognosis Link
@@ 34 MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems Mechanical component
@@
Signals: vibration Diagnosis Link
@@ 35 NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). doi: 10.36001/phmconf.2023.v15i1.3463 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 36 NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Mechanical component
@@
Signals: vibration Prognosis Link
@@ 37 NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: impedance, voltage Prognosis Link
@@ 38 NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: capacity, resistance Prognosis Link
@@ 39 NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Materials
@@
Signals: operating condition, strain, vibration, x-ray Prognosis Link
@@ 40 NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 41 NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: , current, temperature, voltage Prognosis Link
@@ 42 NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 43 NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Manufacturing process
@@
Signals: acoustic emission, current, vibration Prognosis Link
@@ 44 NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 45 NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 46 NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
@@
Signals: current, temperature, voltage Diagnosis Link
@@ 47 NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
@@
Signals: anonymized, operating condition Prognosis Link
@@ 48 NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
@@
Signals: anonymized, operating condition Prognosis Link

### v3

@@ 21 Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
@@
Signals: position, speed, torque
Fault detection Link
@@ 22 Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi
Production system
@@
Signals: unknown
Fault detection Link
@@ 23 Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Signals: flow rate, pressure
Prognosis Link
@@ 24 Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS
Production system
@@
Signals: anonymized
Fault detection Link
@@ 25 Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Signals: dust feed, flow rate, pressure
Prognosis Link
@@ 26 Kaggle - Pump
Kaggle User: UnknownClass
Process technology
@@
Signals: anonymized
Fault detection Link
@@ 27 Kaggle - Sensor Fault Detection
Schneider-Electric, France
Electrical component
@@
Signals: temperature, voltage
Fault detection Link
28 Kaggle - Similar System Data Set for Condition Prognosis
Braig, M.; Zeiler, P. (2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis. IEEE Access, Volume 13, pp. 98485-98503, DOI: 10.1109/ACCESS.2025.3576435
Process technology
Filtration
Signals: flow rate, pressure
Prognosis Link
29 Kaggle - Similar System Data Set for Fault Diagnosis
Braig, M.; Zeiler, P. (2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis. IEEE Access, Volume 13, pp. 98485-98503, DOI: 10.1109/ACCESS.2025.3576435
Mechanical component
Bearing
Signals: vibration
Diagnosis Link
@@ 30 Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
@@
Signals: control parameters
Fault detection Link
31 Karlsruhe Institute of Technology - Ball Screw Drive Surface Defect Dataset
Schlagenhauf, T. (2021): Ball Screw Drive Surface Defect Dataset for Classification. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000133819
Mechanical component
Ball screw drive
Signals: image
Fault detection Link
32 Karlsruhe Institute of Technology - Industrial Machine Tool Element Surface Defect Dataset
Schlagenhauf, T.; Landwehr, M.; Fleischer, J. (2021): Industrial Machine Tool Element Surface Defect Dataset. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000129520
Mechanical component
Ball screw drive
Signals: image
Prognosis Link
@@ 33 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 34 Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58bdhy2df.2
Drive technology
@@
Signals: current, position, temperature
Prognosis Link
@@
Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1
Electrical component
@@
Signals: capacity
Prognosis Link
@@ 36 Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
Drive technology
@@
Signals: pressure, temperature, vibration
Diagnosis Link
@@ 37 Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
Mechanical component
@@
Signals: vibration
Diagnosis Link
@@ 38 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1
Electrical component
@@
Signals: unknown
Prognosis Link
@@ 39 Mendeley - NMC cell 2600 mAh cyclic aging data
Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1
Electrical component
@@
Signals: current, temperature
Prognosis Link
@@ 40 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024). Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435.
Mechanical component
@@
Signals: vibration
Prognosis Link
@@ 41 MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
Mechanical component
@@
Signals: vibration
Diagnosis Link
@@ 42 NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proceedings of the Annual Conference of the PHM Society, 15(1). doi: 10.36001/phmconf.2023.v15i1.3463
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 43 NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Mechanical component
@@
Signals: vibration
Prognosis Link
@@ 44 NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: impedance, voltage
Prognosis Link
@@ 45 NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: capacity, resistance
Prognosis Link
@@ 46 NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Materials
@@
Signals: operating condition, strain, vibration, x-ray
Prognosis Link
@@ 47 NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 48 NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 49 NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 50 NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Manufacturing process
@@
Signals: acoustic emission, current, vibration
Prognosis Link
@@ 51 NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 52 NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 53 NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
@@
Signals: current, temperature, voltage
Diagnosis Link
@@ 54 NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
@@
Signals: anonymized, operating condition
Prognosis Link
@@ 55 NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
@@
Signals: anonymized, operating condition
Prognosis Link